# Improving Vision Transformers by Overlapping Heads in Multi-Head Self-Attention

### v1

@@ I Introduction
Since the introduction of Transformer [1] from language to vision, Vision Transformers have gradually dominated many vision tasks such as image recognition [2][3][4] and object detection [5][6]. Due to the global attention mechanism in Transformers, Vision Transformers could yield better performance in vision tasks when the training data is abundant and the training epochs are enough. Although the convergence rate of Vision Transformers is frequently slower than their Convolutional counterparts, global information communication is one of the key elements for the success of Transformer models.
@@
  • Various variants based on the overlap dimensions are proposed to thoroughly investigate the optimal performance of MOHSA. The proposed MOHSA could be integrated into most Vision Transformer models to enhance their performance with negligible overhead.
@@ II-A Vision Transformers
Vision Transformer [2] utilizes the image patches which are embedded as tokens and a class token as the token inputs for the Transformer encoder to recognize the images. Swin-Transformer [3] reduces the computational cost by limiting the implementation of attention calculation in each window and expands the view of the tokens by shifting the windows. Shuffle Transformer [10] proposes spatial shuffle to exchange information across windows. MSG-Transformer [11] harnesses MSG tokens to represent local windows and the information communication between windows is implemented with MSG tokens. PVT [12] designs Vision Transformer with a hierarchical structure without convolutions by shrinking the feature maps gradually. CaiT [13] and DeepViT [14] investigate Vision Transformers with deeper layers. MobileViT series [15][16][17] and EdgeViTs [18] explore Vision Transformers on mobile-level applications. Some works [19][20][21][22][23] introduce convolutions into the Vision Transformers to take advantage of both convolutions and Transformers. PiT [24] explores the possibility of using individual pixels instead of patches as the tokens for Vision Transformers.
@@ III Approach
Figure 2: Our proposed Multi-Overlapped-Head Self-Attention. MHSA represents the original implementation of Multi-Head Self-Attention with hard division of heads and MOHSA indicates our proposed Multi-Overlapped-Head Self-Attention with soft division of heads. In the original Vision Transformer (left), $Q$, $K$, and $V$ are split for different heads and the attention is computed for each head independently. To exchange the information between heads when the attention is calculated, we propose to overlap $Q$, $K$, $V$ with $Q$, $K$, $V$ in adjacent heads (right). Since overlapped heads would slightly increase the number of dimensions, the projection matrix would project the concatenated heads to the original token dimension.
@@ III-A Multi-Head Self-Attention
Vision Transformer [2] is the Transformer model applied to vision tasks, mainly image classification and recognition. For Vision Transformer, only the encoder is utilized for feature extraction. The images are divided as same-size patches which are embedded as tokens for the Vision Transformer. Layer Normalization [26] is often employed as the normalization before MHSA and FFN. Vision Transformer [2] could be applied to vision tasks and achieve significant performance on vision tasks due to the effective information exchange for attention mechanism. the self-attention mechanism calculates the dot product of each token with all other tokens effectively exchanging the information between them. In vision tasks like image classification and recognition, each patch token has a global view of the image, which is important for each patch to attain the context information.
@@ III-C Overlapping Ratios
In addition to the fixed overlap dimensions for all layers, we also implement some variants of the overlapping ratios by changing the overlap dimensions according to the depths of the layers. In this paper, one layer includes the attention module and the FFN module. The variants for the overlapping ratios are demonstrated in Table I. In Table I, “inc (x layers)” represents the overlap dimension is increased by 1 every x layers, and “dec (x layers)” indicates the overlap dimension is decreased by 1 every x layers, which is a reverse process of ”inc (x layers)” based on the depths of the layers. In addition, “0-indexed” illustrates the overlap dimension starts from 0 for “inc” and ends to 0 for “dec”, and “1-indexed” demonstrates the overlap dimension starts from 1 for “inc” and ends to 1 for “dec”. For instance, inc (2 layers) with “0-indexed” (represented by “inc-0 (2)” in the experiments) for a total of 12 layers has overlap dimensions (0, 0, 1, 1, 2, 2, 3, 3, 4, 4, 5, 5) for layer from 1 to 12, and dec (1 layer) with “1-indexed” (represented by “dec-1 (1)” in the experiments) for total 12 layers has overlap dimensions (12, 11, 10, 9, 8, 7, 6, 5, 4, 3, 2, 1) for layer from 1 to 12.
@@ IV-A CIFAR-10
The ablation study for ViT-Tiny and ViT-Small on CIFAR-10 is demonstrated in Table III. In Table III, “fixed 1” represents the overlap dimension is 1 for all layers, and “fixed half” indicates the overlap dimension is half of the head dimension for all layers. “inc” indicates the overlap dimension is increased with the depths of the layers increasing and “dec” represents the overlap dimension is decreased with the layers going deeper. The two numbers following “inc” or “dec”, for instance, “inc-0 (3)” illustrates the overlap dimension is increased by 1 every 3 layers and it is “0-indexed”. More details about the variants of overlapping ratios can be found in Sec. III-C.
The best accuracy for ViT-Tiny is 85.65 for variant “inc-0 (6)” and the best accuracy for ViT-Small is 87.30 for variant “inc-0 (1)”. For ViT-Tiny, “fixed half” demonstrates higher accuracy than “fixed 1”. For ViT-Small, “fixed half” has lower accuracy than “fixed 1”. For a fixed paradigm of overlap dimension, more ratios of overlapping cannot guarantee higher accuracy.
@@ IV-B CIFAR-100
Table V demonstrates the experiments of ViT-Tiny and ViT-Small on CIFAR-100. The best result for ViT-Tiny is 63.01 with “inc-0 (2)” and for ViT-small is 64.97 with “dec-1 (3)”. Both results improve the models by more than 1% than the original models. Similar to the results on CIFAR-10, “fixed half” has higher accuracy for ViT-Tiny, and “fixed 1” illustrates better performance for ViT-Small. In addition, “fixed half” does not demonstrate effectiveness for ViT-Small.
Table VI illustrates the performance of CaiT on CIFAR-100. For CaiT-xxs12, using 1 overlap dimension for all layers could boost the accuracy by nearly 2%. For CaiT-xxs24, the accuracy is significantly enhanced by 5% with “inc-1 (3)”. Additionally, “inc-1 (3)” is much better than using half dimension of the head as the overlap dimension which utilizes more parameters and FLOPs.
@@ IV-C Tiny-ImageNet
The experimental results on Tiny-ImageNet are demonstrated in Table VII. The accuracy of ViT-Tiny could be enhanced by 1.61% utilizing “inc-1 (2)”. ViT-Small is boosted by 1.27% with “inc-1 (1)”. Moreover, using half of the head dimension as the overlap dimension is better than using 1 as the overlap dimension for all layers for ViT-Tiny, while it is not the case for ViT-Small.
For CaiT-xxs12 and CaiT-xxs24, “fixed 1” that utilizes fixed overlap dimension 1 for all layers could significantly improve the accuracy by 2.39% and 3.58%, respectively. Furthermore, the accuracy is enormously enhanced by 7.41% for CaiT-xxs24 with “inc-1 (1)”. The effectiveness of our method is significant on CaiT models and we can ignore the slightly increased parameters and computations.
For Swin-Tiny, “fixed 1” paradigm could greatly enhance the accuracy by 1.23%, and the performance could be further boosted by a large margin with “dec-1 (1)”. The effectiveness of overlapped heads is also demonstrated on the Transformer model with window-based hierarchical architecture.
@@ IV-D ImageNet
By training and testing on ImageNet, we could investigate the effectiveness of our proposed approach on large datasets. The experimental results on ImageNet are demonstrated in Table VIII. Almost all models equipped with our proposed MOHSA have significant improvements on ImageNet.
@@
Figure 4: The accuracy comparison during the training process. The accuracy comparison between the original method and our proposed approach for CaiT-xxs24 on val or test set of CIFAR-100, Tiny-ImageNet and ImageNet is illustrated from left to right. The blue curve demonstrates our approach with the best performance and the red curve indicates the original method.
@@ IV-E Analysis
In the experiments, we utilize different variants of our proposed MOHSA on various Vision Transformer models and datasets to illustrate the effectiveness of MOHSA. Overall, the enhancement of the models on various datasets is remarkable with such insignificant overhead. For different variants of our proposed method, the results manifest some differences. For a fixed paradigm that the overlap dimension is the same for all layers, the performance could be enhanced by only 1 overlap dimension in most cases and increasing the overlap dimension for fixed mode cannot guarantee better results. Additionally, the variants of varying overlap dimensions based on the depths of the layers demonstrate superior performance than the fixed paradigm on various models and datasets in most cases. Compared to the fixed paradigm with a high overlap dimension, varying overlap dimensions with the depths of the layers could save the number of parameters and computational costs.
Fig. 4 manifests the accuracy of CaiT-xxs24 on the val or test set with the training proceeding between the original models and our proposed models. “inc-1 (3)”, “inc-1 (1)” and “inc-0 (1)” are selected as ours for comparison for CIFAR-100, Tiny-ImageNet, and ImageNet, respectively. The accuracy between the original models and our models demonstrates almost no difference at the early stage of training, while diverges significantly with the training proceeding.
@@ V Conclusion
In this paper, we have proposed a simple yet effective module MOHSA, to improve the original MHSA in Vision Transformers by overlapping the heads, allowing for information exchange during attention calculation in each head. To the best of our knowledge, this is the first work to propose overlapping heads and achieve significant enhancement across various datasets for different Vision Transformer models. We hope our work will inspire the community to further explore the structure of Vision Transformers.

### v2

@@ Abstract
Index Terms:
Vision Transformers, Attention Mechanism, Multi-Head Self-Attention
@@ I Introduction
Since the introduction of Transformer [1] from language to vision, Vision Transformers have gradually dominated many vision tasks such as image recognition [2][3][4] and object detection [5][6][7][8]. Due to the global attention mechanism in Transformers, Vision Transformers could yield better performance in vision tasks when the training data is abundant and the training epochs are enough. Although the convergence rate of Vision Transformers is frequently slower than their Convolutional counterparts, global information communication is one of the key elements for the success of Transformer models.
@@
  • Various variants based on the overlap dimensions are proposed to thoroughly investigate the optimal performance of MOHSA. The proposed MOHSA could be integrated into Vision Transformer models to enhance their performance with negligible overhead.
@@ II-A Vision Transformers
Vision Transformer [2] utilizes the image patches which are embedded as tokens and a class token as the token inputs for the Transformer encoder to recognize the images. Swin-Transformer [3] reduces the computational cost by limiting the implementation of attention calculation in each window and expands the view of the tokens by shifting the windows. Shuffle Transformer [12] proposes spatial shuffle to exchange information across windows. MSG-Transformer [13] harnesses MSG tokens to represent local windows and the information communication between windows is implemented with MSG tokens. PVT [14] designs a Vision Transformer with a hierarchical structure without convolutions by shrinking the feature maps gradually. CaiT [15] and DeepViT [16] investigate Vision Transformers with deeper layers. A multi-layer dense attention decoder was proposed in [17] for segmentation. MobileViT series [18][19][20] and EdgeViTs [21] explore Vision Transformers on mobile-level applications. Some works [22][23][24][25][26] introduce convolutions into the Vision Transformers to take advantage of both convolutions and Transformers. PiT [27] explores the possibility of using individual pixels instead of patches as the tokens for Vision Transformers. ViTAR [28] adjusts Vision Transformers on various image resolutions. PartialFormer [29] proposes partial attention for more efficient Vision Transformers on various vision tasks.
@@ II-B Attention Heads Interaction
Figure 2: The Transformer encoder for a typical Transformer model. The Transformer encoder is exploited in Vision Transformer for image classification. $N$ indicates the number of layers for the Transformer encoder.
@@ III Approach
Figure 3: Our proposed Multi-Overlapped-Head Self-Attention. MHSA represents the original implementation of Multi-Head Self-Attention with hard division of heads and MOHSA indicates our proposed Multi-Overlapped-Head Self-Attention with soft division of heads. In the original Vision Transformer (left), $Q$, $K$, and $V$ are split for different heads and the attention is computed for each head independently. To exchange the information between heads when the attention is calculated, we propose to overlap $Q$, $K$, $V$ with $Q$, $K$, and $V$ in adjacent heads (right). Since overlapped heads would slightly increase the number of dimensions, the projection matrix would project the concatenated heads to the original token dimension.
@@ III-A Multi-Head Self-Attention
Vision Transformer [2] is the Transformer model applied to vision tasks, mainly image classification and recognition. For Vision Transformer, only the encoder is utilized for feature extraction, as demonstrated in Fig. 2. The images are divided as same-size patches which are embedded as tokens for the Vision Transformer. Layer Normalization [31] is often employed as the normalization before MHSA and FFN. Vision Transformer [2] could be applied to vision tasks and achieve significant performance on vision tasks due to the effective information exchange for attention mechanism. the self-attention mechanism calculates the dot product of each token with all other tokens effectively exchanging the information between them. In vision tasks like image classification and recognition, each patch token has a global view of the image, which is important for each patch to attain the context information.
@@ III-C Overlapping Ratios
In addition to the fixed overlap dimensions for all layers, we also implement some variants of the overlapping ratios by changing the overlap dimensions according to the depths of the layers. In this paper, one layer includes the attention module and the FFN module, as illustrated in Fig. 2 which represents one layer of the Transformer encoder. The variants for the overlapping ratios are demonstrated in Table I. In Table I, “inc (x layers)” represents the overlap dimension is increased by 1 every x layers, and “dec (x layers)” indicates the overlap dimension is decreased by 1 every x layers, which is a reverse process of ”inc (x layers)” based on the depths of the layers. In addition, “0-indexed” illustrates the overlap dimension starts from 0 for “inc” and ends at 0 for “dec”, and “1-indexed” demonstrates the overlap dimension starts from 1 for “inc” and ends at 1 for “dec”. For instance, inc (2 layers) with “0-indexed” (represented by “inc-0 (2)” in the experiments) for a total of 12 layers has overlap dimensions (0, 0, 1, 1, 2, 2, 3, 3, 4, 4, 5, 5) for layer from 1 to 12, and dec (1 layer) with “1-indexed” (represented by “dec-1 (1)” in the experiments) for total 12 layers has overlap dimensions (12, 11, 10, 9, 8, 7, 6, 5, 4, 3, 2, 1) for layer from 1 to 12.
@@ IV-A CIFAR-10
CIFAR-10 [9] includes 50k training images and 10k testing images. Among all 60k images, there are 10 classes and each class has 6k images. The ablation study for ViT-Tiny and ViT-Small on CIFAR-10 is demonstrated in Table III. In Table III, “fixed 1” represents the overlap dimension is 1 for all layers, and “fixed half” indicates the overlap dimension is half of the head dimension for all layers. “inc” indicates the overlap dimension is increased with the depths of the layers increasing and “dec” represents the overlap dimension is decreased with the layers going deeper. The two numbers following “inc” or “dec”, for instance, “inc-0 (3)” illustrates the overlap dimension is increased by 1 every 3 layers and it is “0-indexed”. More details about the variants of overlapping ratios can be found in Sec. III-C.
The best accuracy for ViT-Tiny is 85.65 for variant “inc-0 (6)” which increases the overlap dimension by 1 every 6 layers with “0-indexed” and the best accuracy for ViT-Small is 87.30 for variant “inc-0 (1)” which increases the overlap dimension by 1 every 1 layer with “0-indexed”. For ViT-Tiny, “fixed half” demonstrates higher accuracy than “fixed 1”. For ViT-Small, “fixed half” has lower accuracy than “fixed 1”.
@@ IV-B CIFAR-100
CIFAR-100 [9] includes 50k training images and 10k testing images. There are 100 classes and each class has 600 images. Table V demonstrates the experiments of ViT-Tiny and ViT-Small on CIFAR-100. The best result for ViT-Tiny is 63.01 with “inc-0 (2)” which increases the overlap dimension by 1 every 2 layers with “0-indexed” and for ViT-small is 64.97 with “dec-1 (3)” which decreases the overlap dimension by 1 every 3 layers with “1-indexed”. Both results improve the models by more than 1% than the original models. Similar to the results on CIFAR-10, “fixed half” has higher accuracy for ViT-Tiny, and “fixed 1” illustrates better performance for ViT-Small. In addition, “fixed half” does not demonstrate effectiveness for ViT-Small.
Table VI illustrates the performance of CaiT on CIFAR-100. For CaiT-xxs12, using 1 overlap dimension for all layers could boost the accuracy by nearly 2%. For CaiT-xxs24, the accuracy is significantly enhanced by 5% with “inc-1 (3)” which increases the overlap dimension by 1 every 3 layers with “1-indexed”. Additionally, “inc-1 (3)” is much better than using half dimension of the head as the overlap dimension which utilizes more parameters and FLOPs.
@@ IV-C Tiny-ImageNet
Tiny-ImageNet [10] includes 100k training images and 10k validation images. There are 200 classes in the dataset. Each class has 500 images for training and 50 images for validation. The experimental results on Tiny-ImageNet are demonstrated in Table VII. The accuracy of ViT-Tiny could be enhanced by 1.61% utilizing “inc-1 (2)” that increases the overlap dimension by 1 every 2 layers with “1-indexed”. ViT-Small is boosted by 1.27% with “inc-1 (1)” that increases the overlap dimension by 1 every 1 layer with “1-indexed”. Moreover, using half of the head dimension as the overlap dimension is better than using 1 as the overlap dimension for all layers for ViT-Tiny, while it is not the case for ViT-Small.
For CaiT-xxs12 and CaiT-xxs24, “fixed 1” that utilizes fixed overlap dimension 1 for all layers could significantly improve the accuracy by 2.39% and 3.58%, respectively. Furthermore, the accuracy is enormously enhanced by 7.41% for CaiT-xxs24 with “inc-1 (1)” which increases the overlap dimension by 1 every 1 layer with “1-indexed”. The effectiveness of our method is significant on CaiT models and we can ignore the slightly increased parameters and computations.
For Swin-Tiny, “fixed 1” paradigm could greatly enhance the accuracy by 1.23%, and the performance could be further boosted by a large margin with “dec-1 (1)” that decreases the overlap dimension by 1 every 1 layer with “1-indexed”. The effectiveness of overlapped heads is also demonstrated on the Transformer model with window-based hierarchical architecture.
Figure 5: The accuracy comparison during the training process. The accuracy comparison between the original method and our proposed approach for CaiT-xxs24 on val or test set of CIFAR-100, Tiny-ImageNet, and ImageNet is illustrated from left to right. The blue curve demonstrates our approach with the best performance and the red curve indicates the original method.
@@ IV-D ImageNet
ImageNet-1k [11] includes 1k classes with more than 1 million training images and 50k validation images. By training and testing on ImageNet, we could investigate the effectiveness of our proposed approach on large datasets. The experimental results on ImageNet are demonstrated in Table VIII. “inc-0 (1)” and “inc-1 (1)” indicate the overlap dimension is increased by 1 every 1 layer with “0-indexed” and “1-indexed”, respectively. “dec-0 (1)” and “dec-1 (1)” represent the overlap dimension is decreased by 1 every 1 layer with “0-indexed” and “1-indexed”, respectively. Almost all models equipped with our proposed MOHSA have significant improvements on ImageNet.
@@ IV-E Analysis
In the experiments, we utilize different variants of our proposed MOHSA on various Vision Transformer models and datasets to illustrate the effectiveness of MOHSA. Overall, the enhancement of the models on various datasets is remarkable with such insignificant overhead. Our proposed approach might be exploited as a plug-and-play method for Vision Transformer models and might enhance the performance of those models with minimum extra cost. For different variants of our proposed method, the results manifest some differences. For a fixed paradigm that the overlap dimension is the same for all layers, the performance could be enhanced by only 1 overlap dimension in most cases and increasing the overlap dimension for fixed mode cannot guarantee better results. Additionally, the variants of varying overlap dimensions based on the depths of the layers demonstrate superior performance than the fixed paradigm on various models and datasets in most cases. Compared to the fixed paradigm with a high overlap dimension, varying overlap dimensions with the depths of the layers could save the number of parameters and computational costs.
To illustrate the accuracy during training, Fig. 5 manifests the accuracy of CaiT-xxs24 on the val or test set with the training proceeding between the original models and our proposed models. “inc-1 (3)”, “inc-1 (1)” and “inc-0 (1)” are selected as ours for comparison for CIFAR-100, Tiny-ImageNet, and ImageNet, respectively. The accuracy between the original models and our models demonstrates almost no difference at the early stage of training, while diverges significantly with the training proceeding.
@@ V Conclusion
In this paper, we have proposed a simple yet effective module MOHSA, to improve the original MHSA in Vision Transformers by overlapping the heads, allowing for information exchange during attention calculation in each head. Extensive evaluations and comparisons with the state-of-the-arts on multiple datasets have demonstrated the superior performance of the proposed approach. To the best of our knowledge, this is the first work to propose overlapping heads and achieve significant enhancement across various datasets for different Vision Transformer models. We hope our work will inspire the community to further explore the structure of Vision Transformers.